OpenStudy (anonymous):

Help with operations with complex numbers? 7i*3i(-8-6i).

OpenStudy (anonymous):

168+126i

OpenStudy (masumanwar):

since i*i =-1 7i*3i*8=168 and 7i*3i*6i=126i